OpenStudy (kevin):

xyz = 2^6 and (2 log x)( 2 log yz) + (2log y)(2 log z) = 10 with x,y,z >= 0. Then sqrt(2 log^2 x + 2 log^2 y + 2 log^2 z) = ....

OpenStudy (kevin):

\[xyz = 10^6\] and \[(\log _{2} x) (\log _{2} yz) + (\log _{2} y) (\log _{2} z) = 10 \] with x,y,z >= 0 Then \[\sqrt{\log_{2}^{2}x + \log _{2}^{2}y + \log_{2}^{2}z} = ....\]

OpenStudy (kevin):

@agent0smith @3mar

OpenStudy (3mar):

Where you got such problems, Kevin?

OpenStudy (kevin):

all. lol

OpenStudy (3mar):

you need x and y and z. right?

OpenStudy (kevin):

The answer is 4, maybe you can explain to me about how to get that

OpenStudy (kevin):

no... find sqrt of bla bla bla as I wrote above

OpenStudy (3mar):

I see

OpenStudy (3mar):

I will do my best In sha' Allah

OpenStudy (3mar):

Let's begin.

OpenStudy (kevin):

thx :D

OpenStudy (3mar):

You are welcome!

OpenStudy (3mar):

Do you have Skype Id?

OpenStudy (3mar):

It will be more easier there!

OpenStudy (kevin):

@Directrix @IrishBoy123 @Astrophysics @SapphireMoon @sweetburger @jackthegreatest

OpenStudy (kevin):

Sorry, but I don't have :(

OpenStudy (kevin):

Hi @Will.H

OpenStudy (3mar):

for \[xyz=10^6\] take the log to the base 2 for both sides \[\log_{2} (xyz)=\log_{2}(10^6) \]

OpenStudy (3mar):

then multiply both sided to 2 \[2\log_2(xyz)=2\log_210^6\]

OpenStudy (3mar):

As you know the feature of the exponent of the logarithmic function: \[alogx=\log(x^2)\]

OpenStudy (3mar):

sorry \[a*\log(x)=\log(x^a)\]

OpenStudy (3mar):

Do you follow?

OpenStudy (kevin):

yup

OpenStudy (kevin):

just go on :)

OpenStudy (3mar):

then it would take the form: \[\log_2(xyz)^2=2\log_210^6\] then distribute the power 2 to x, y and z \[\log_2(x^2.y^2.z^2)=2\log_210^6\]

OpenStudy (3mar):

then separete what inside the log() by plus sign as follows: \[\log_2(x^2)+\log_2(y^2)+\log_2(z^2)=2\log_210^6\] then take the square root for both sides: \[\sqrt{\log_2(x^2)+\log_2(y^2)+\log_2(z^2)}=\sqrt{2\log_210^6}=number\] That's it. Done!

OpenStudy (3mar):

Hope that helps

OpenStudy (3mar):

got it or stick?

OpenStudy (kevin):

actually the question is sqrt of 2 log 2 x + blablabla.... not sqrt of 2 log (x^2) ..... Is it just me that I don't understand? Thx anyway, I've figured out how is it after following your first method on the post above. I got stick at that before.

OpenStudy (kevin):

It's like this...